OpenStudy (anonymous):

Given a geometric sequence in the table below, create the explicit formula and list any restrictions to the domain. n an 1 −4 2 20 3 −100

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

What's the common ratio?

OpenStudy (shamil98):

\[\huge x_n = ar^{n-1}\]

OpenStudy (shamil98):

a = first term r = ratio n = number of terms

OpenStudy (anonymous):

pls teach me how to figure out the common ratio @kc_kennylau

OpenStudy (kc_kennylau):

\[\Large r=a_2\div a_1\]

OpenStudy (shamil98):

so what is \[20 \div -4\]?

OpenStudy (solomonzelman):

\[r=a_{n} \div a_{n-1}\]

OpenStudy (kc_kennylau):

@shamil98 @woohoo sorry i got to go now :)

OpenStudy (anonymous):

ok so r=-100/-4

OpenStudy (shamil98):

anyways you have: n = 3 a = -4 r = -5 So using the formula i gave you, just plug them in

OpenStudy (shamil98):

a_2 = 20 a_1 = -4 20 / - 4 = -5

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

400?

OpenStudy (shamil98):

\[\Huge x_n = ar^{n-1} \] \[\large \text{You have the values for a,r, and n.}\] \[\large \text {r = -5 , n = 3 , a = -4}\] \[\large \text{So,now, just plug them into the formula.}\]

OpenStudy (shamil98):

Well, actually, it says create the formula for it, so leave n just as n.

OpenStudy (anonymous):

an = −5(−4)n − 1 where n ≥ 1

OpenStudy (shamil98):

Correct, i assume you mean to the power of n-1. \[\large a_n = -5(-4)^{n-1}\]

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

isnt this one wrong? do you remember?

OpenStudy (kc_kennylau):

Why is this wrong?

OpenStudy (anonymous):

It was wrong when I tried.

OpenStudy (solomonzelman):

wait what ? what was wrong ?

OpenStudy (anonymous):

an = −5(−4)n − 1 where n ≥ 1 the answer

OpenStudy (anonymous):

@SolomonZelman an = −5(−4)n − 1 where n ≥ 1 an = −4(−5)n − 1 where n ≥ 1 an = −4(5)n − 1 where n ≥ −4 an = 5(−4)n − 1 where n ≥ −4 those are the answers wouldn't it be the last one

OpenStudy (solomonzelman):

answers for what question ?

OpenStudy (anonymous):

this one

OpenStudy (solomonzelman):

these are the options you posted just now, and the question ?

OpenStudy (anonymous):

Given a geometric sequence in the table below, create the explicit formula and list any restrictions to the domain. n an 1 −4 2 20 3 −100 woohoo's original question.

OpenStudy (solomonzelman):

I don't even know what to say... I am just going to say that the question says "Given a geometric sequence in the table below \(\normalsize\color{red}{ \bf in ~the~ table~below }\)."

OpenStudy (anonymous):

an = 5(−4)n − 1 where n ≥ −4 was wrong...hm...I have no idea what the right answer was.

OpenStudy (anonymous):

It's the second option that xoalybabyxo posted.